Solve the system by the elimination method
5x+3y=-11
7x-2y=13
The solution is ?
The solution is supposed to be an ordered pair of integers or fractions for any numbers in the expression.
I have tried to get the solution over and over to no success could someone please help me? Thank You so much.

Answer by Alan3354(69443)

5x+3y=-11
7x-2y=13
----------
10x + 6y = -22 1st eqn x 2
21x - 6y = 39 2nd eqn x 3
------------- Add
31x = 17
x = 17/31
------------
5*(17/31) + 3y = -11
3y = -11 - 85/31 = -426/31
y = -142/31
--> (17/31,-142/31)